The vapour pressure of a moist gas at 100 degree celcius is 750 mm and aqueous tension at temperature is 10 mm then vapour pressure of the dry gas.

Grade:12th pass

1 Answers

Kiran Satheesh
6 years ago
From the given data we haveP(moist) = P(dry) + aqueous tension750 = P(dry) + 10P(dry) = 740 mm Hence the answer.